# Dual Data Alignment Makes AI-Generated Image Detector Easier Generalizable

### v2

@@ 2 Related Works
Table 1: Comparing on dataset alignment.
@@ 3.1 Motivation and Analysis
Without additional supervision, detectors rely solely on the training set to learn the concept of ’real vs. synthetic’. When these two classes systematically differ in attributes like image size, style, or semantics, the model may mistakenly treat these non-causal disparities as defining characteristics of ’real’ versus ’fake.’ In particular, some of these discrepancies are particularly evident and easy for models to detect. However, genuine real-synthetic artifacts are subtle and often imperceptible, making them harder for detectors to learn the true unbiased features.
@@
Figure 5: Illustration of DDA pipeline. Left: VAE-reconstructed images differ from real images in the intensity of high-frequency components. Right: DDA aligns images by first fusing high-frequency information from real images into the VAE-reconstructed images to align them in the frequency domain. Then, DDA uses pixel-level mixup of real and frequency-aligned images to further align them in the pixel domain.
@@ 5 Conclusion
While our method demonstrates superior performance on both benchmark datasets and real-world scenarios, it may struggle to detect images with strong post-processing artifacts, as shown in the results on the Chameleon dataset. In future work, we plan to extend DDA to synthetic videos to enhance generalization in AI-generated video detection.

### v3

@@ 1 Introduction
Table 1: Comparing on dataset alignment.
Figure 3: Visual illustration of how dataset bias affects decision boundaries in synthetic image detection. Left: Detectors trained on biased datasets—where synthetic images (e.g., Syn.1–3) differ from real images in format, content, or resolution—tend to learn spurious decision boundaries. These boundaries may misclassify samples that deviate from the biased training distribution. Right: When synthetic images are carefully aligned with real images across multiple domains (e.g., format, content, resolution), the model learns a tighter, more generalizable decision boundary. This alignment enables better detection of previously unseen synthetic images by covering a wider distributional range.
@@ 2 Related Works
3 Methodology
3.1 Motivation and Analysis
@@
In the absence of additional supervision, detectors must rely exclusively on the training set to learn the concept of ’real’ versus ’synthetic’. When these two classes differ systematically in non-causal attributes—such as image size, compression format, or semantic content—the model may incorrectly learn to associate these irrelevant features with authenticity. These spurious signals are often more salient than subtle, genuine artifacts that actually distinguish real from synthetic images, making it more difficult for the model to learn truly generalizable, causally relevant features. Figure 3 illustrates how such biases constrain the learned decision boundary, as shown in the left three panels. In contrast, the rightmost panel shows that when synthetic images are carefully aligned with real images across key domains, the model can learn a tighter, more transferable decision boundary, improving generalizability to unseen data.
@@
Figure 6: DDA pipeline. Left: VAE-reconstructed images differ from real ones in the intensity of high-frequency components. Right: DDA fuses high-frequency information from real images into the VAE-reconstructed images to align them in the frequency domain. Then, DDA uses pixel-level mixup of real and frequency-aligned images to further align them in the pixel domain.
@@ 5 Conclusion
While our method shows strong performance on both benchmarks and real-world data, it remains sensitive to heavily post-processed images, as seen in the Chameleon dataset. Interestingly, we also observed that even photos taken by smartphones may exhibit synthetic-like artifacts—likely due to the AI-based enhancement features embedded in modern smartphone camera pipelines. In future work, we plan to: (1) clarify the boundary between real and synthetic images in post-processed content, and (2) extend DDA to video-based detection tasks to support generalization in AI-generated video detection.